The length of a rectangle is 25 more inches more than its width. If a 2 inch square is cut from each corner and the sides are folded up, the resulting open box has a volume of 372 cubic inches. Find the width of the original rectangle.
Answer by checkley77(12844)

L=W+25
2(L-2*2)(W-2*2)=372
2(W+25-4)(W-4)=372
2(W+21)(W-4)=372
W^2+21W-4W-84=372/2
W^2+17W-84=186
W^2+17W-84-186=0
W^2+17W-270=0
(W+27)(W-10)=0
W-10=0
W=10 ANS. FOR THE WIDTH.
L=10+25
L=35 ANS. FOR THE LENGTH.
PROOF:2(35-4)(10-4)=372
2*31*6=372
372=372